# An Efficient Row-Based Sparse Fine-Tuning

### v1

@@ 6.1 Main Results of LLM
We apply our SPruFT method to fine-tune Llama2-7B and Llama3-8B, comparing the results with those obtained through LoRA and its variants. We select the magnitude of the neuron vector as the importance metric due to its low memory requirements, simplicity, and widely tested effectiveness. In contrast, gradient-based metrics like Taylor and Hessian are as memory-intensive as full LLM fine-tuning. While Wanda (Sun et al., 2024) offers a memory-efficient metric for pruning LLMs, it still requires one epoch of data forwarding and significantly more memory than inference to compute the input vector’s norm [^6] We encountered an OOM error when using Wanda’s official implementation. When pruning a neural network, each layer computes Wanda and is pruned immediately, gradually reducing the model size. However, in the fine-tuning process, the model size remains unchanged. Additionally, storing the norm values for computing importance scores further increases memory consumption, making memory cost significantly higher than when using Wanda for pruning.. For epochs choosing, we opt for 5 epochs to balance computational resources and performance.
Table 3 demonstrates the exceptional memory efficiency of our approach [^7] Also refer to Table 12 in Appendix C.6, even with $r=128$, our method’s memory usage remains significantly lower than that of LoRA with $r=16$. while achieving comparable accuracy. As shown, the accuracies of fine-tuned models remain similar across most PEFT methods, while memory usage varies significantly. VeRA, despite having significantly fewer trainable parameters, shows noticeably lower accuracy. Notably, our approach consistently requires substantially less memory than all other PEFT methods listed in the table.
Model, ft setting mem #param Avg
Llama2(7B), LoRA, $r=16$ 21.64GB 40.0M(0.59%) 60.55
Llama2(7B), FA-LoRA, $r=32$ 16.56GB 46.4M(0.69%) 60.71
Llama2(7B), LoRA, $r=32$ 22.21GB 80.0M(1.19%) 60.92
Llama2(7B), FA-LoRA, $r=64$ 17.25GB 92.8M(1.38%) 60.85
Llama2(7B), SPruFT, $r=32$ 15.57GB 36.4M(0.54%) 60.52
Llama2(7B), FA-SPruFT, $r=64$ 14.48GB 39.3M(0.58%) 60.82
Llama2(7B), SPruFT, $r=64$ 16.20GB 72.9M(1.08%) 60.67
Llama2(7B), FA-SPruFT, $r=128$ 15.21GB 78.6M(1.17%) 60.87
Llama3(8B), LoRA, $r=16$ 28.86GB 41.9M(0.52%) 64.76
Llama3(8B), FA-LoRA, $r=32$ 23.89GB 56.6M(0.71%) 64.81
Llama3(8B), LoRA, $r=32$ 29.37GB 83.9M(1.04%) 64.76
Llama3(8B), FA-LoRA, $r=64$ 24.55GB 113.2M(1.41%) 64.77
Llama3(8B), SPruFT, $r=32$ 22.62GB 39.8M(0.50%) 64.21
Llama3(8B), FA-SPruFT, $r=64$ 21.62GB 46.1M(0.57%) 64.56
Llama3(8B), SPruFT, $r=64$ 23.23GB 79.7M(0.99%) 64.51
Llama3(8B), FA-SPruFT, $r=128$ 22.41GB 92.3M(1.15%) 64.26
@@
Table 4: Comparing fine-tuning all linear layers with fine-tuning only the MLP layers. FA indicates that we freeze attention layers, but not including MLP layers followed by attention blocks. Full table please refer to Table 14 in Appendix C.7.
@@
We then demonstrate that strategically assigning trainable parameters saves more memory than merely reducing them, without compromising accuracy. We compare fine-tuning all linear layers and fine-tuning only the MLP layers, with results shown in Table 4. The former requires more memory for the same number of trainable parameters, as distributing trainable parameters across the model increases the need for caching intermediate values. Recalling the results from Table 1 in Section 3, freezing attention blocks provides significantly greater memory savings than reducing LoRA’s trainable parameters. Additionally, accuracy remains nearly unchanged. Given that Llama models have been pre-trained on extensive datasets, their attention blocks likely already capture crucial patterns for token interactions. Our results suggest that freezing these attention blocks maintains performance nearly equivalent to fine-tuning all layers.
@@ 6.2 Importance Metrics
Settings Avg
Llama2(7B) Llama3(8B)
$r=128$, random 60.25 64.00
$r=128$, $\ell^{2}$ 60.86 64.50
$r=128$, ZOTaylor 60.94 64.67
$r=128$, FA, random 60.35 63.95
$r=128$, FA, $\ell^{2}$ 60.87 64.26
$r=128$, FA, ZOTaylor 60.87 64.38
Table 6: Importance evaluation for Llama2 and Llama3. Random row selection serves as the baseline here. We also present the results of freezing-attention blocks in this table (FA). Full table with different ranks please refer to Table 13 in Appendix C.6.
@@
Similarly, in the cases of Llama2 and Llama3, our findings suggest that ZOTaylor provides a slight performance boost for fine-tuned models. This improvement is likely due to ZOTaylor’s ability to capture richer data information, whereas magnitude-based importance tends to focus more on identifying generally important neurons. However, the observed performance gain remains modest, potentially due to the variance of the estimates, as discussed in Section 4.2.2. Beyond these observations, another interesting finding is that fine-tuned models with random row selection perform similarly to VeRA, likely suggesting that this accuracy level could serve as a baseline for other fine-tuning approaches.
@@ 7 Conclusions and Future Work
We propose a parameter-efficient fine-tuning (PEFT) framework that integrates various techniques and importance metrics from model compression research to enhance sparse fine-tuning (SFT). Using our method, we can fine-tune LLMs and vision transformers using significantly less computation resources than the popular LoRA (Low-Rank Adaptation) technique, while achieving similar accuracy. We also explore the effects of using different importance metrics. There are several future directions: (1) For importance metrics, while Quantile-Mean Taylor shows slight improvements, these gains are relatively minor compared to the standard Taylor metric in some cases of DeiT and ViT. We may wish to explore better metrics for classification tasks with a large number of labels. (2) Developing memory-efficient importance metrics for LLMs is another future direction. While Zeroth-Order Taylor is effective for incorporating data-specific information without requiring large memory, the large variance of estimate is a challenge. Although we reduce the variance effectively by increasing the number of estimations, exploring a simple method to reduce variance without increasing estimation times is essential for further advancements in this field. (3) Our results show that fine-tuning a small number of neurons can significantly improve model performance on downstream tasks. This observation naturally raises the question: do the selected neurons play a distinctive role in specific tasks? This question is related to the explainability of neural networks, which is an extensive area of research. It will be interesting to understand if (and how) individual neurons chosen for fine-tuning contribute to the new task.
@@ Appendix A Importance Metrics
Taylor importance is the Taylor expansion of the difference between the loss of the model with and without the target neuron:
${\bm{\eta}}_{i}$ $=L(\mathcal{D},\acute{F}_{\bm{c}_{i}})-L(\mathcal{D},F)$
$\approx-{\mathbf{w}}^{\top}\nabla_{{\mathbf{w}}}L(\mathcal{D},F)+\frac{1}{2}{\mathbf{w}}^{\top}\nabla_{{\mathbf{w}}}^{2}L(\mathcal{D},F){\mathbf{w}}$
$\overset{(*)}{\approx}\frac{1}{2}{\mathbf{w}}^{\top}\nabla_{{\mathbf{w}}}^{2}L(\mathcal{D},F){\mathbf{w}}$
$\overset{(**)}{\approx}\frac{1}{2}(G{\mathbf{w}})^{\top}(G{\mathbf{w}}),$
where $G=\nabla_{{\mathbf{w}}}L(\mathcal{D},F)$. (**) is from the result of Fisher information (Rissanen, 1996):
@@
${\bm{\eta}}_{i}\approx(G{\mathbf{w}})^{\top}(G{\mathbf{w}})=\sum_{j}(\frac{1}{|\mathcal{D}|}\sum_{{\mathbf{x}}\in\mathcal{D}}\frac{\partial L({\mathbf{x}},F)}{\partial w_{j}}w_{j})^{2}\approx\sum_{j}|\frac{1}{|\mathcal{D}|}\sum_{{\mathbf{x}}\in\mathcal{D}}\frac{\partial L({\mathbf{x}},F)}{\partial w_{j}}w_{j}|,$
where ${\mathbf{w}}=(w_{1},\ldots,w_{j},\ldots)$.
Magnitude importance is the $\ell_{2}$-norm of the neuron vector computed as $\sqrt{\sum_{j}w^{2}_{j}}$.

### v2

@@ 6.1 Main Results of LLM
We apply our SPruFT method to fine-tune Llama2-7B and Llama3-8B, comparing the results with those obtained through LoRA and its variants. We select the magnitude of the neuron vector as the importance metric due to its low memory requirements, simplicity, and widely tested effectiveness. In contrast, gradient-based metrics like Taylor and Hessian are as memory-intensive as full LLM fine-tuning. While Wanda (Sun et al., 2024) and AWQ (Lin et al., 2024) offer two activation-based metrics to evaluate neuron importance, for pruning LLMs, they require one epoch of data forwarding and significantly more memory than inference to compute the activation-based importance [^7] We encountered an OOM error when using Wanda’s official implementation. When pruning a neural network, each layer computes activation vector’s norm and is pruned immediately, gradually reducing the model size. However, in the fine-tuning process, the model size remains unchanged. Additionally, storing the activation values for computing importance scores further increases memory consumption, making memory cost significantly higher than when using activation for pruning..
@@
Table 3: Comparing fine-tuning all linear layers with fine-tuning only the FFN on commonsense reasoning tasks. indicates that we freeze the layers for queue, key, and value projection. Full table please refer to Table 17 in Appendix D.7.
Models Llama2(7B) Llama3(8B)
Setting mem #param Common(Avg) mem #param Common(Avg)
LoRA, $r=16$ 21.64GB 40.0M(0.59%) 72.48 28.86GB 41.9M(0.52%) 78.95
LoRA, $r=32$ 17.95GB 54.8M(0.81%) 72.90 25.28GB 65.0M(0.81%) 78.35
LoRA, $r=32$ 22.21GB 80.0M(1.19%) 72.03 29.37GB 83.9M(1.04%) 79.50
LoRA, $r=64$ 18.81GB 109.6M(1.63%) 73.00 26.04GB 130.0M(1.62%) 77.73
SPruFT, $r=32$ 15.57GB 36.4M(0.54%) 72.25 22.62GB 39.8M(0.50%) 78.53
SPruFT, $r=64$ 14.67GB 47.7M(0.71%) 72.45 21.81GB 54.5M(0.68%) 79.05
SPruFT, $r=64$ 16.20GB 72.9M(1.08%) 72.65 23.23GB 79.7M(0.99%) 77.88
SPruFT, $r=128$ 15.58GB 95.4M(1.42%) 73.98 22.71GB 109.1M(1.36%) 80.45
Table 2 demonstrates the exceptional memory efficiency of our approach [^8] Also refer to Table 14 in Appendix D.6, even with $r=128$, our method’s memory usage remains significantly lower than that of LoRA with $r=16$. while achieving comparable accuracy. As shown, the accuracies of fine-tuned models remain similar across most PEFT methods, while memory usage varies significantly. VeRA, despite having significantly fewer trainable parameters, shows noticeably lower accuracy. Notably, our approach consistently requires substantially less memory than all other PEFT methods listed in the table.
Table 4: MT-Bench results of fine-tuning full precision Llama2 and Llama3 on Alpaca-GPT4. Bold indicates the best result on the same task.
Model, ft setting Coding Extraction Humanities Math Reasoning Roleplay Stem Writing Avg
Llama2(7B), pretrained 1.00 2.63 1.65 1.50 1.59 1.65 1.70 1.55 1.66
LoRA, $r=64$ 1.50 2.59 5.21 2.43 3.27 5.40 4.59 3.84 3.60
VeRA, $r=64$ 1.47 2.75 2.30 1.42 3.06 2.55 3.06 2.10 2.34
DoRA, $r=64$ 1.13 2.55 4.22 1.93 3.50 4.15 4.35 4.35 3.27
RoSA, $r=32,d=1.2\%$ 1.73 3.35 5.00 2.64 4.90 4.80 3.94 3.84 3.78
SPruFT, $r=128$ 1.82 2.55 4.80 2.08 4.07 4.79 4.50 3.53 3.52
LoRA, $r=64$ 2.36 1.89 4.89 1.92 3.36 4.30 4.05 3.84 3.33
VeRA, $r=64$ 1.47 2.50 2.15 1.39 3.06 2.90 2.32 1.85 2.20
DoRA, $r=64$ 1.10 1.75 4.25 2.46 3.62 4.89 4.21 3.58 3.23
RoSA, $r=32,d=1.2\%$ 2.00 3.11 5.60 2.50 4.43 4.50 4.11 3.74 3.75
SPruFT, $r=128$ 1.08 2.10 4.60 1.15 3.80 4.22 4.33 3.42 3.09
Llama3(8B), pretrained 3.56 5.56 2.95 1.32 2.00 2.84 3.56 3.75 3.19
LoRA, $r=64$ 2.88 3.88 5.05 4.33 4.25 5.89 5.72 5.22 4.65
VeRA, $r=64$ 3.18 5.28 4.05 1.00 2.39 3.55 4.72 3.00 3.40
DoRA, $r=64$ 3.82 4.85 6.40 5.00 4.00 6.21 5.88 4.65 5.10
RoSA, $r=32,d=1.2\%$ 3.50 7.31 6.30 4.31 4.58 5.75 7.19 5.30 5.53
SPruFT, $r=128$ 3.88 5.11 6.11 3.83 4.21 5.35 6.44 5.40 5.04
LoRA, $r=64$ 3.75 4.78 6.15 4.42 4.75 5.15 6.06 5.95 5.12
VeRA, $r=64$ 3.33 5.32 4.35 1.00 1.76 3.50 4.25 4.00 3.44
DoRA, $r=64$ 3.90 5.56 6.75 4.42 3.82 6.50 6.88 5.05 5.36
RoSA, $r=32,d=1.2\%$ 3.63 5.88 6.05 4.43 2.92 5.32 6.87 5.30 5.05
SPruFT, $r=128$ 4.75 4.78 5.16 3.67 3.31 6.25 6.06 5.00 4.87
@@
We then demonstrate that strategically assigning trainable parameters saves more memory than merely reducing them, without compromising accuracy on commonsense reasoning tasks. Section 3 highlights the importance of the computation graph in memory consumption. Strategically assigning trainable parameters can be an effective solution. For instance, Shi et al. (2024) suggest that layers such as output, down, up, and gate projections contribute more significantly than queue, key, and value projections, which leads to the strategy of freezing queue, key, and value projections. We compare fine-tuning all linear layers and fine-tuning only the FFN and output-project layer, with results shown in Table 3. The former requires more memory for the same number of trainable parameters, as distributing trainable parameters across the model increases the need for caching intermediate values. Table 3 shows an exceptional memory saving by freezing some layers in attention blocks. In addition, the accuracy remains nearly unchanged. Given that Llama models have been pre-trained on extensive datasets, their attention layers likely already capture crucial patterns for token interactions.
We further evaluate models’ instruction-following abilities using MT-Bench. As shown in Table 4, SPruFT with magnitude-based importance falls short of other methods. However, incorporating ZOTaylor significantly improves performance: Table 6 (Section 6.2) shows that SPruFT with ZOTaylor outperforms most approaches on instruction-following tasks, ranking second in all cases. Only RoSA performs better overall, but it requires substantially more memory than our method.
@@ 6.2 Importance Metrics
Table 6: Importance evaluation for Llama2 and Llama3 on commonsense reasoning tasks, GSM8k, and MT-Bench. We also present the results of freezing queue-, key-, and value-projection in this table (). Full table with different ranks for commonsense tasks please refer to Table 15 and Table 16 in Appendix D.6.
Model Llama2(7B) Llama3(8B)
Settings Common(Avg) GSM8K MT-Bench(Avg) Common(Avg) GSM8k MT-Bench(Avg)
$r=128$, random 69.80 22.52 3.45 76.08 42.53 4.40
$r=128$, $\ell^{2}$ 74.53 22.90 3.52 80.55 46.10 5.04
$r=128$, ZOTaylor 74.35 23.50 3.62 81.28 55.12 5.34
$r=128^{{\ddagger}}$, random 72.08 21.08 3.60 77.20 39.50 5.05
$r=128^{{\ddagger}}$, $\ell^{2}$ 73.98 19.48 3.09 80.45 42.53 4.87
$r=128^{{\ddagger}}$, ZOTaylor 74.30 24.34 3.66 81.80 46.02 5.21
@@
Similarly, in the cases of Llama2 and Llama3, our findings suggest that ZOTaylor provides a slight performance boost for fine-tuned models. This improvement is likely due to ZOTaylor’s ability to capture richer data information, whereas magnitude-based importance tends to focus more on identifying generally important neurons. However, the observed performance gain remains modest, potentially due to the variance of the estimates, as discussed in Section 4.3.2. Beyond these observations, an interesting finding is that models fine-tuned with random row selection significantly outperform VeRA, likely suggesting that overly aggressive parameter reduction can substantially degrade performance.
6.3 Results of Quantized Models
Table 7: Main results of fine-tuning NF4-quantized LLaMA-2 and LLaMA-3. QSPruFT denotes our approach with $\mathrm{\Delta}{\mathbf{W}}$ initialized randomly, while QSPruFT uses $\mathrm{\Delta}{\mathbf{W}}$ initialized with the corresponding rows from the quantization residual ${\mathbf{W}}^{res}$.
Model, ft setting BoolQ PIQA SIQA HS WG ARC-c ARC-e OBQA Avg GSM8k
Llama2(7B), pretrained 37.00 37.00 19.20 20.00 4.40 15.40 17.80 20.80 21.48 0.00
pretrained-NF4 58.00 40.00 29.00 15.40 4.80 14.00 16.20 25.40 25.35 0.00
QLoRA, $r=64$ 75.20 74.80 66.80 73.80 63.60 63.40 77.60 65.40 70.08 18.88
QDoRA, $r=64$ 77.40 72.00 68.40 81.60 63.20 64.60 80.20 70.60 72.25 25.63
LoftQ, $r=64$ 73.80 73.20 72.60 81.60 64.20 67.60 82.00 71.60 73.33 20.55
QPiSSA, $r=64$ 79.40 72.40 68.60 77.40 65.00 67.40 78.60 69.80 72.33 21.08
QSPruFT, $r=128$ 78.00 72.60 69.00 79.40 63.20 67.20 79.60 74.80 72.98 20.32
QSPruFT, $r=128$ 82.00 71.80 69.20 83.20 65.40 67.20 79.20 72.40 73.80 22.14
Llama3(8B), pretrained 58.80 41.60 38.00 10.20 11.20 55.20 63.00 27.40 38.18 0.00
pretrained-NF4 52.00 37.80 11.60 8.40 0.20 48.80 53.80 16.20 28.60 0.00
QLoRA, $r=64$ 77.40 81.60 72.80 87.60 69.60 75.80 90.20 78.60 79.20 41.17
QDoRA, $r=64$ 79.00 79.20 68.20 82.20 66.40 75.80 87.40 77.20 76.93 46.17
LoftQ, $r=64$ 87.00 82.80 69.40 90.20 62.40 74.80 91.00 79.60 79.65 44.96
QPiSSA, $r=64$ 87.60 82.00 71.00 82.60 71.80 77.20 89.60 77.20 79.88 45.87
QSPruFT, $r=128$ 89.80 79.40 65.40 89.80 68.40 76.20 91.80 78.00 79.85 47.76
QSPruFT, $r=128$ 90.40 82.40 68.60 89.40 69.20 76.80 90.80 76.40 80.50 49.13
Tables 7 and 8 present the results of fine-tuning NF4-quantized models using various PEFT methods [^9] We omit the results of QPiSSA in Table 8 because its training loss was not reasonably low in our experiments. While the original paper (Meng et al., 2024) suggests disabling dropout to accelerate convergence, we observed that doing so leads to overfitting in all tasks. Instead, we set the dropout rate to $0.1$ (consistent with LoRA’s setting), which yielded better performance on commonsense reasoning tasks and GSM8k. However, for AlpacaGPT4, the training loss of QPiSSA remained above $2$, whereas all other methods achieved losses below $0.1$. One possible explanation for the underperformance of QPiSSA is its slight modification of the frozen parameter matrix ${\mathbf{W}}^{\text{NF4}}$. Although the matrix $\hat{{\mathbf{W}}}$ in QPiSSA before training is similar to those in other methods, QPiSSA changes the frozen matrix ${\mathbf{W}}^{\text{NF4}}$ slightly, which may require more training data to achieve good performance.. In these quantized experiments, we use magnitude-based importance metric to save processing time, as quantized parameters increase latency during training and inference. The results show that our method, with $\mathrm{\Delta}{\mathbf{W}}$ initialized from ${\mathbf{W}}^{res}$, consistently achieves strong performance across tasks. These findings support that SPruFT is a simple, effective, and memory-efficient fine-tuning strategy.
Table 8: Main results of fine-tuning NF4-quantized LLaMA-2 and LLaMA-3. QSPruFT denotes our approach with $\mathrm{\Delta}{\mathbf{W}}$ initialized randomly, while QSPruFT uses $\mathrm{\Delta}{\mathbf{W}}$ initialized with the corresponding rows from the quantization residual ${\mathbf{W}}^{res}$.
Model, ft setting Coding Extraction Humanities Math Reasoning Roleplay Stem Writing Avg
Llama2(7B), pretrained 1.00 2.63 1.65 1.50 1.59 1.65 1.70 1.55 1.66
pretrained-NF4 1.00 2.84 1.60 1.19 1.67 1.85 1.80 1.40 1.67
QLoRA, $r=64$ 1.00 3.05 4.74 1.79 3.23 5.15 4.69 3.55 3.40
QDoRA, $r=64$ 1.46 3.78 5.42 1.92 3.63 4.63 4.63 4.05 3.69
LoftQ, $r=64$ 1.46 3.90 5.55 2.45 4.69 4.60 4.37 3.89 3.87
QSPruFT, $r=128$ 0.88 3.75 5.65 2.29 2.77 4.58 3.82 4.40 3.52
QSPruFT, $r=128$ 1.92 4.70 5.45 1.93 4.57 5.20 4.78 4.60 4.14
Llama3(8B), pretrained 3.56 5.56 2.95 1.32 2.00 2.84 3.56 3.75 3.19
pretrained-NF4 1.57 4.32 2.35 1.75 2.50 2.80 3.11 2.21 2.58
QLoRA, $r=64$ 4.43 5.41 4.74 3.00 4.79 5.65 4.74 4.68 4.68
QDoRA, $r=64$ 3.56 6.26 6.00 4.43 3.83 5.95 5.28 5.06 5.05
LoftQ, $r=64$ 4.30 7.00 6.60 5.55 3.73 5.90 5.60 6.30 5.62
QSPruFT, $r=128$ 2.43 3.74 4.26 2.90 3.40 3.53 3.89 3.95 3.51
QSPruFT, $r=128$ 4.50 5.41 7.06 4.00 4.62 5.42 5.79 5.80 5.32
@@ 7 Conclusions and Future Work
We propose a parameter-efficient fine-tuning (PEFT) framework that integrates various techniques and importance metrics from model compression research to enhance sparse fine-tuning (SpFT). Using our method, we can fine-tune LLMs and vision transformers using significantly less computation resources than the popular LoRA (Low-Rank Adaptation) technique, while achieving similar accuracy. We also explore the effects of using different importance metrics. There are several future directions: (1) For importance metrics, while Quantile-Mean Taylor shows slight improvements, these gains are relatively minor compared to the standard Taylor metric in some cases of DeiT and ViT. We may wish to explore better metrics for classification tasks with a large number of labels. (2) Developing memory-efficient importance metrics for LLMs is another future direction. While Zeroth-Order Taylor is effective for incorporating data-specific information without requiring large memory, the large variance of estimate is a challenge. Although we reduce the variance effectively by increasing the number of estimations, exploring a simple method to reduce variance without increasing estimation times is essential for further advancements in this field. (3) Our results show that fine-tuning a small number of neurons can significantly improve model performance on downstream tasks. This observation naturally raises the question: do the selected neurons play a distinctive role in specific tasks? This question is related to the explainability of neural networks, which is an extensive area of research. It will be interesting to understand if (and how) individual neurons chosen for fine-tuning contribute to the new task.
@@ Appendix A Importance Metrics
Taylor importance is the Taylor expansion of the difference between the loss of the model with and without the target neuron: (we denote ${\bm{\theta}}^{(i)}$ by ${\mathbf{w}}$ here for convenience.)
${\bm{\eta}}_{i}$ $=L(\mathcal{D},\acute{F}_{i})-L(\mathcal{D},F)$
$\approx-{\mathbf{w}}^{\top}\nabla_{{\mathbf{w}}}L(\mathcal{D},F)+\frac{1}{2}{\mathbf{w}}^{\top}\nabla_{{\mathbf{w}}}^{2}L(\mathcal{D},F){\mathbf{w}}+\mathcal{O}(\nabla_{{\mathbf{w}}}^{3}L(\mathcal{D},F))$
$\overset{(*)}{\approx}\frac{1}{2}{\mathbf{w}}^{\top}\nabla_{{\mathbf{w}}}^{2}L(\mathcal{D},F){\mathbf{w}}+\mathcal{O}(\nabla_{{\mathbf{w}}}^{3}L(\mathcal{D},F))$
$\overset{(**)}{\approx}\frac{1}{2}(G{\mathbf{w}})^{\top}(G{\mathbf{w}})+\mathcal{O}(\nabla_{{\mathbf{w}}}^{3}L(\mathcal{D},F)),$
where $G=\nabla_{{\mathbf{w}}}L(\mathcal{D},F)$. (**) is from the result of Fisher information Rissanen (1996):
@@
${\bm{\eta}}_{i}\approx(G{\mathbf{w}})^{\top}(G{\mathbf{w}})$ $=\sum_{j}(\frac{1}{|\mathcal{D}|}\sum_{{\mathbf{x}}\in\mathcal{D}}\frac{\partial L({\mathbf{x}},F)}{\partial{\mathbf{w}}_{j}}{\mathbf{w}}_{j})^{2}$ (14)
$\approx\sum_{j}|\frac{1}{|\mathcal{D}|}\sum_{{\mathbf{x}}\in\mathcal{D}}\frac{\partial L({\mathbf{x}},F)}{\partial{\mathbf{w}}_{j}}{\mathbf{w}}_{j}|^{2},$ (15)
where ${\mathbf{w}}=({\mathbf{w}}_{1},\ldots,{\mathbf{w}}_{j},\ldots)$. Thus, people often compute $\sum_{j}|\frac{1}{|\mathcal{D}|}\sum_{{\mathbf{x}}\in\mathcal{D}}\frac{\partial L({\mathbf{x}},F)}{\partial{\mathbf{w}}_{j}}{\mathbf{w}}_{j}|$ without taking square to evaluate the importance score. In this study, we describe this as Taylor importance and consider eq 14 as hessian importance.
Magnitude importance is the $\ell_{2}$-norm of the neuron vector computed as $\sqrt{\sum_{j}{\mathbf{w}}^{2}_{j}}$.
Appendix B Missing Proofs